OpenStudy (anonymous):

I need help with this question! I don't understand how to find the answer please help! The sum of three consecutive integers is 99. What is the lowest of the three integers? a) 17 b) 28 c) 30 d) 32

OpenStudy (anonymous):

If you work up with it, like take each number and add them to see if its right

OpenStudy (anonymous):

17+18+19 = 28+29+30= 30+31+32= 32+33+34=

OpenStudy (anonymous):

well consecutive means any numbers that come right after one another, like 2,3,4 or 45,46,47, or like 33,34,35, lets try 17,18, and 19 they add up to 54, so thats not it b and c dont work either so its d

OpenStudy (anonymous):

I believe

OpenStudy (anonymous):

lax us you need to let her do it so she understands unless they just ask for the answer

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Thank you @jsouza and @moudown828 I didn't know what consecutive meant and I tried it by myself! Thanks :)

OpenStudy (anonymous):

youre welcome XD if you need help with anything let me know

OpenStudy (amistre64):

even if they just ask the answer ... we still do not just give them an answer out right